Last year, Dan had
$30,000
to invest. He invested some of it in an account that paid
5%
simple interest per year, and he invested the rest in an account that paid
10%
simple interest per year. After one year, he received a total of
$1950
in interest. How much did he invest in each account?

Answer by ewatrrr(24785)

Some at 5%, rest at 10%
Use .05 and .10 for the Arithmetic
Let x be the amount at 10%
.10x + .05(30000 - x) = 1950
.05x = 1950 - 1500
x = 450/.05
x = $9000, the amount at 10%. 30000-9000 = $21,000 at 5%
CHECK...
$900 + $1050 = $1950